OpenStudy (anonymous):

2x^2 - 8 How would I factor this? I mean, I CAN get it factored, but I spend far too much time hunting for an answer. I know there are more optimal methods out there of factoring polynomials like this, so I figured I'd ask here.

hartnn (hartnn):

take 2 common and then use \(\huge a^2-b^2=(a+b)(a-b)\)

OpenStudy (anonymous):

both terms are multiples of 2 then use difference of squares

OpenStudy (anonymous):

I generally start this kind with (2x )(x ) But I have a hard time with the ones that have anything other than 1 in the leading term.

hartnn (hartnn):

2(x^2-4) = 2(x^2-2^2) i hope u can take it further.

OpenStudy (anonymous):

I would apply what hartnn suggested. It's the right approach.

OpenStudy (anonymous):

@halimabegum you cannot solve it, its an expression not an equation

OpenStudy (anonymous):

2x^2 -8 = 2 ( x^2 - 4)

OpenStudy (anonymous):

2[(x^2)-(2^2)] =2((x-2)(x+2)) difference of two squares or 2(x-2)(x+2)

OpenStudy (anonymous):

So I can just factor out the 2? \[2(x^{2}-4)\] and then use the difference of square from there? 2 ((x-2)(x+2)) When would I be able to tell that this method (factoring something out and then using the difference of squares) is applicable? Unless I would look for something to factor out right off the bat... would I?

OpenStudy (anonymous):

2x^2-8 2( x^2 -4) 2 ( x + 2) ( x- 2)

hartnn (hartnn):

take some other example of your own....with leading co-efficient not 1,. and try this method.

OpenStudy (anonymous):

Okay, @hartnn \[-5x ^{2}+5\] I would factor out the 5: \[5(-x ^{2}+1)\] 5(-x+1)(x+1) ?

hartnn (hartnn):

nopes., take -5 out.

hartnn (hartnn):

make leading co=efficient = 1, not -1

OpenStudy (anonymous):

What would happen to the 5? -1?

OpenStudy (anonymous):

Okay, so I would convert it to 5x^2 - 5?

hartnn (hartnn):

-1(5x^2-5) = -5 (x^2-1)

OpenStudy (anonymous):

-5(x+1)(x-1)

hartnn (hartnn):

yup.

OpenStudy (anonymous):

Sweet, thanks bro!

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

@melbel what you did earlier 5(1-x^2) = 5(-x+1)(x+1) <---- this is fine too. it's 5(1+x)(1-x)